# SoK: On Finding Common Ground in Loss Landscapes Using Deep Model Merging Techniques

### v1

@@ VII-A Loss Landscape Macrophenomena
Mode Determinism: The final linearly connected objective basin to which deep learning models converge is determined early in the training process. This assertion is supported by experiments on linear connectivity [47] and by visualizations of model clustering in parameter space [82]. SGD has been observed to occur in two phases: an initial period of instability, followed by a more stable, linear trajectory [47, 120, 121, 122, 123, 124]. Fort et al. [62] show that this behavior corresponds to the expected dynamics of the Neural Tangent Kernel. Weight aggregation techniques implicitly use mode determinism; otherwise, they could not merge models using linear combinations of their parameters. Instead, we would expect barriers of high loss between models that were independently fine-tuned. Linearly connected models also demonstrate linearity in their features—that is, linear interpolations between their parameters correspond to linear interpolations between their feature maps [39, 111]. This result implies that the nature of the representations that models will learn is also determined early in the training process. Damian et al. [121] explain the stability in the later part of the training trajectory for neural networks by the tendency of SGD to implicitly regularize the sharpness of the loss landscape. They demonstrate the explicit presence of such a regularization term in the Taylor expansion of the SGD update equation. This result suggests that the observed stability in the training trajectory is a property of SGD itself.
@@
Mode Directedness: Within a given objective basin, we can meaningfully associate directions in parameter space with task-specific behavior. This “directedness” is best illustrated by the success of methods such as Task Arithmetic [85]. Follow-up work, such as TIES-merging [88] and DARE [89], demonstrate that removing redundancies and low-magnitude differences between task vectors improves merging performance. This implies that just a few principal directions can effectively account for task-specific behavior. Gueta et al. [82] demonstrate the clustering behavior of language models trained on similar tasks. They observe that models trained on the same dataset are tightly clustered in parameter space, whereas models trained on the same underlying task with different datasets form looser clusters. This notion of hierarchically organized task-specific areas in parameter space offers an explanation for the success of several standard routines in the contemporary deep learning landscape, including the pre-train fine-tune training paradigm and meta-learning procedures such as Model-Agnostic Meta-Learning [125]. These procedures may move parameters into regions of space where effective parameter configurations for more specialized sub-tasks are abundant. While the existence of meaningful task-specific regions in space has been proposed previously to motivate meta-learning approaches, the merging literature justifies these conjectures with empirical evidence showing that this kind of organization is present at multiple scales.
@@
Mode Connectivity: Even solutions found in distinct objective modes can be connected by low-loss paths that are comprised of equally performant solutions. Moreover, this connectivity between solutions in separate modes is commonplace, as is LMC, i.e., the presence of linear paths of low loss between known solutions in objective space. There may even be entire manifolds of low loss connecting equally performant solutions [72]. An accumulation of evidence suggests that models trained on similar tasks, even with different initializations, can be connected by paths of low loss if one accounts for the inherent permutation invariance of neural networks by appropriately aligning their parameters after training [16, 19, 20, 47, 51]. Additionally, Li et al. [57] showcase how constraining training of a network to a subspace or pruning prior to training reduces the number of inherent permutation symmetries and leads to LMC between solutions without further post-training alignment. This result implies that constraining neuron misalignment from arising can enable the discovery of connected paths between modes more easily.
@@
Several theoretical results suggest that the loss landscape of overparameterized models is dominated by subspaces of permutation-equivalent objective basins [55]. This observation suggests that model representations are generally quite similar—even under different forms of SGD noise or initializations, they tend to converge to similar solutions. This similarity may be even more prevalent as models are further over-parameterized, offering an explanation for why merging methods are often more successful as model size increases.
VII-B Implications for Model Training
@@
The direct comparison of model representations in merging studies allows us to derive information about the effects of model architecture and pre-training objectives on learned representations. Here, we outline some of these findings.
@@
Model Size: Neuron alignment procedures struggle as model width decreases. Ito et al. [59] investigate this phenomenon by studying the distribution of singular values in the weight matrices of source models. They conjecture that the effectiveness of neuron alignment techniques is predicated on their ability to align preferentially the singular directions corresponding to dominant singular values. They find that the proportion of dominant singular values decreases substantially as model width increases, suggesting that wider models are easier to align. Nguyen et al. [117] show empirical results suggesting that over-parameterized models propagate dominant principal components of their hidden representations at each layer. These results support the idea that a few dominant feature directions arise in over-parameterized models.
@@
Aghajanyan et al. [15] measure the intrinsic dimensionality of language models and similarly conclude that LLMs form more compressed representations as they scale in size. These works suggest that larger models, despite their increased expressiveness, have fewer dominant directions in feature space. Some work has also attempted to clarify why over-parameterized models may display more amicable optimization behavior than smaller models. Huang et al. [66] and Li et al. [14] visualize the loss landscapes of models at various sizes and find that overparameterized models often converge to large, flat basins correlated with enhanced generalization capability [64, 66, 67, 68, 69, 126, 127] as opposed to small sharp minima. They explain this tendency by observing that such basins occupy far more volume in objective space as the dimensionality of a model increases. The increased volume of flat basins, coupled with the results presented by Simsek et al. [55] indicating that the number of permutation-equivalent basins increases exponentially as models grow in size, may explain the relative ease with which large models train: the volume of objective space corresponding to good solutions grows exponentially with model size.
@@
Figure 6: Mode Directedness: Optimal parameter settings for related tasks lie clustered together in parameter space; clusters become looser as tasks diverge in similarity. This suggests a hierarchical organization of the objective landscape with respect to semantically similar tasks.
@@
Pre-training Objective: Models with different pre-training objectives display different sensitivities to various merging techniques [21, 26]. For example, transformer models in the T5 family [128] perform well even with only simple averaging; meanwhile, DeBERTa models [129] consistently degrade in performance when merged, even when using state-of-the-art techniques [21]. This discrepancy points to an interplay between the pre-training objective and the success of model merging procedures. One possible explanation for this difference is the effect of pre-training objective on training dynamics. Some evidence suggests a strong link between a model’s pre-training objective and the final objective basin in which it lies—for example, contrastive pre-training has been observed to coerce models to lie in flatter objective basins associated with better generalization and robustness guarantees [38]. Certain pre-training procedures seem to predispose models toward representations that display linear connectivity and linear relationships between activations [21, 111]. Understanding this interaction is important for interpretability work that seeks to isolate task-specific behavior [130, 131, 132] or to disentangle and identify key directions in latent space [133]. Some work has explored the overlap between pre-training objective, loss landscape geometry, and adversarial robustness [38, 134]. Further work might consider the causal effects of pre-training objective on loss landscape geometry and relate it to resulting model robustness and performance in downstream tasks.
VIII Discussion
Model merging is closely tied to model interpretability. Techniques in interpretability research, such as activation steering, can control language style, reduce model toxicity, and edit factual associations post-training [135, 136]. Such methods involve altering model behavior via linear changes to the activation space [137]. Existing work on task arithmetic [85], TIES-merging [88], and AdaMerge [97] demonstrates the viability of this approach to model steering by showcasing the linear correspondence between changes in model activations and changes in model weights [39, 111]. Model merging techniques provide valuable insights into the mutable nature of hidden representations and their connection to model parameters. The linearity of feature representations implies that cheap interventions on model weights can replicate such procedures—circumventing the need to apply activation steering methods that require extensive tuning [135, 136].
Interpretability studies should focus on groups of models. Model merging literature highlights that machine learning practitioners have the tools to localize task-specific behaviors to specific model components (e.g., layers, neurons, weights). These localization tools can enable researchers to interpret how a model learns information. Conventionally, interpretability studies seek to understand the underlying mechanics of individual models [29, 96, 130, 131, 137, 138]; however, this paradigm is limiting.
@@ VIII Discussion
We propose that studying groups of models, as done in the model merging literature, would more easily enable researchers to make substantial claims about how common behaviors (e.g., memorization) arise across families of models. Interpretability work can draw upon insights from the model merging literature, for instance: (i) models trained on similar tasks can be aligned to find common subsets of parameters for a given task [19, 20, 101]; (ii) models trained on different randomly chosen subsets of data or different hyperparameters when merged can help identify parts of the models that are relevant to the overall task invariant to the specific data or hyperparameters that was used to train the model [78, 79, 83, 101]; and (iii) models with similar underlying training distributions cluster together in parameter space [82].
For example, to study the behavior of memorization in LLMs, we argue that it is easier to glean insights into how a model learns that behavior by studying how (and if) it arises in a group of models with variations in their training recipes (as opposed to studying a single model exhibiting memorization). Methods such as ZipIt! [101] showcase how we can distinguish parameters associated with differences in behavior between models. We can also find new interventions on model parameters using notions of loss landscape geometry. For example, Barbulescu and Triantafillou [30] demonstrate how to use Task Arithmetic to unlearn memorized data. We can imagine similar interventions that also consider the mode-directedness of the objective landscape—perhaps identifying whole regions of parameter space that are more prone to degenerate behavior like memorization, and developing train-time mitigations to prevent their development.
The fundamental unit of analysis in neural networks. The discussion in this paper focuses predominantly on merging weights. However, much analysis of neural networks is done in activation space. What the correct unit of analysis is in neural networks is an open question. Neurons are the most natural choice for analysis due to their simplicity. However, rigorous examinations show that natural language explanations for neuron activation patterns systematically display poor precision and recall while also failing to behave as expected under causal interventions [139]. Elhage et al. [140] propose that neural networks represent concepts in “superposition.” That is, rather than having meaningful concepts axis-aligned in their hidden states, neural networks pack many almost orthogonal concepts into the high dimensional space defined by the neurons. This idea has led researchers to focus on analysis in linear directions [141], polytopes [142], and subspaces [143] in the activation space. Dictionary learning [144] and sparse autoencoders [133, 145] identify meaningful directions in the hidden states of neural networks in an unsupervised fashion—both methods learn compressed representations of hidden states by producing sparse reconstructions of them. Ideally, these representations would correspond to interpretable concepts. However, Chaudhary and Geiger [146] find that the features learned by sparse autoencoders perform worse than neuron-level features under causal interventions on a benchmark dataset.
Geiger et al. [95] propose Distributed Alignment Search, a supervised method that learns a rotation matrix that rotates the hidden states of a neural network so that the trained features are axis-aligned. Intuitively, this shows that meaningful features correspond to directions in the hidden state, but these features are not automatically aligned to the axes defined by the neurons. Each of these methods relies on the intuition that the structure of what a neural network has learned is encoded in the geometry of the activation space. Most activation steering work alters activations [29, 137] and relies on the intuition that directions in activation space are the fundamental unit of analysis of neural networks. The model merging literature suggests that directions in weight space may also be appropriate for model steering and analysis.
@@
Advancements in model security depend on a rich understanding of loss landscape geometry. Mode determinism and mode directedness are implicitly or explicitly applied in various attacks on deep learning models. In black-box scenarios, it is common to train local substitute models of a similar architecture on training data for the same underlying task to develop adversarial samples [147, 148]. This assumption of similarity in learned representation and behavior is encapsulated by the concept of mode directedness and is supported by the model merging and loss landscape literature. Many attacks, such as data poisoning and membership inference, can be aided by training similar models on similar data distributions [9]. Mode determinism implies that such attacks may be especially effective when deployed against freely available pre-trained checkpoints, as fine-tuned models will likely be even closer in their representations [47, 82].
@@
The study of model security is increasingly dependent on an understanding of loss landscape geometry [149, 40]. For example, the success of gradient obfuscation against gradient-based adversaries has been linked to how gradient obfuscation alters the local loss landscape geometry—making it difficult to ascend for adversaries. Conversely, this insight led to the development of a countermeasure that enables adversaries to effectively mitigate defenses based on gradient obfuscation by navigating an altered loss geometry [150]. Train-time defenses against adversarial examples frequently involve navigating towards robust local minima via the addition of an adversarial robustness term—one can even thwart subsequent fine-tuning success in restricted domains by appropriate engineering of the loss function [134]. This has been observed to bias models toward large, flat loss basins. Indeed, a strong correlation has been established between adversarial robustness and local loss curvature, indicating that solutions in large, flat basins are more robust to adversaries [149].
Furthermore, the locally convex, globally connected geometry of objective landscapes offers a framework from which defenses may arise. Gradient obfuscation and other schemes that sample model parameters from a distribution implicitly use local convexity in objective basins by sampling predictions from nearby models. Some works have even explicitly exploited mode connectivity to learn paths between vulnerable solutions that contain more robust models—outperforming fine-tuning in terms of model robustness and accuracy [149]. This underscores the advancements that a deep understanding of loss landscape geometry may yield in model security.
IX Conclusion & Future Work

### v2

@@ 6.1.2. Mode Determinism:
The final linearly connected objective basin to which deep learning models converge is determined early in the training process. In particular, models that share a significant portion of their training trajectory (e.g., models fine-tuned from the same pre-trained checkpoint) will converge to solutions in the same objective basin (see Fig. 6). Frankle et al. (40) illustrate the tendency of models to become stable to sources of SGD noise (e.g., batch order, data augmentation) early in training, even observing that at a certain point in the training trajectory, models are determined to lie in the same linearly connected region. This assertion is corroborated by Gueta et al. (44), who observe that fine-tuned language model checkpoints cluster loosely around their pre-trained ancestors, even when trained on semantically dissimilar tasks. SGD has been observed to occur in two phases: an initial period of instability, followed by a more stable, linear trajectory (40; 39; 22; 37; 86; 21). Fort et al. (36) show that this initially chaotic, eventually stable optimization behavior corresponds to the expected dynamics of the Neural Tangent Kernel—observing that within just 2-3 epochs, the final linearly connected region that a solution will lie in is already determined. Weight aggregation techniques implicitly use mode determinism; otherwise, they could not merge models using linear combinations of their parameters. Instead, we would expect barriers of high loss between models that were independently fine-tuned. Linearly connected models also demonstrate linearity in their features—that is, linear interpolations between their parameters correspond to linear interpolations between their feature maps (142; 140). This result implies that the nature of the representations that models will learn is also determined early in the training process. Damian et al. (22) explain the stability in the later part of the training trajectory for neural networks by the tendency of SGD to implicitly regularize the sharpness of the loss landscape. They demonstrate the explicit presence of such a regularization term in the Taylor expansion of the SGD update equation. This result suggests that the observed stability in the training trajectory is a property of SGD itself.
6.1.3. Mode Directedness:
@@
Within a given objective basin, we can meaningfully associate directions in parameter space with task-specific behavior, which is depicted in Fig. 7. This “directedness” is best illustrated by the success of methods such as Task Arithmetic (58), which explicitly construct task vectors that associate task-specific behavior with particular directions in parameter space. Follow-up work, such as TIES-merging (130) and DARE (134), demonstrate that removing redundancies and low-magnitude differences between task vectors improves merging performance. This implies that just a few principal directions in parameter space can effectively account for task-specific behavior. Gueta et al. (44) demonstrate the clustering behavior of language models trained on similar tasks. They observe that models trained on the same dataset are tightly clustered in parameter space, whereas models trained on the same underlying task with different datasets form looser clusters. This notion of hierarchically organized task-specific areas in parameter space offers an explanation for the success of several standard routines in the contemporary deep learning landscape, including the pre-train fine-tune training paradigm and meta-learning procedures such as Model-Agnostic Meta-Learning (35). These procedures may move parameters into regions of space where effective parameter configurations for more specialized sub-tasks are abundant. While the existence of meaningful task-specific regions in parameter space has been proposed previously to motivate meta-learning approaches, the model merging literature justifies these conjectures with empirical evidence showing that this kind of organization is present at multiple scales.
6.1.4. Mode Connectivity:
@@
Even solutions found in distinct objective modes can be connected by low-loss paths that are comprised of equally performant solutions. Moreover, this connectivity between solutions in separate modes is commonplace, as is LMC, i.e., the presence of linear paths of low loss between known solutions in objective space. There may even be entire manifolds of low loss connecting equally performant solutions (9). An accumulation of evidence suggests that models trained on similar tasks, even with different initializations, can be connected by paths of low loss if one accounts for the inherent permutation invariance of neural networks by appropriately aligning their parameters after training (30; 3; 107; 40; 115). Additionally, Li et al. (81) showcase how constraining training of a network to a subspace or pruning prior to training reduces the number of inherent permutation symmetries and leads to LMC between solutions without further post-training alignment. This result implies that constraining neuron misalignment from arising can enable the discovery of connected paths between modes more easily.
@@
Simsek et al. (106) present several theoretical results suggesting that subspaces of permutation-equivalent objective basins corresponding to global minimizers dominate the loss landscape of heavily overparameterized models. This result suggests that vast overparameterization can lead to a more convex objective landscape. Under a quasiconvex objective landscape consisting of many permutation equivalent minima, model representations may generally be quite similar—even under different forms of SGD noise or initializations, they tend to converge to similar solutions modulo a permutation symmetry. This similarity would be even more prevalent as models are further overparameterized, offering an explanation for why merging methods are often more successful as model size increases.
6.2. Implications for Model Training
@@
The direct comparison of model representations in merging studies allows us to derive information about the effects of model architecture and pre-training objectives on the development of learned representations. Here, we outline some of these findings.
6.2.1. Model Size:
@@
Neuron alignment procedures struggle as model width decreases. Ito et al. (61) investigate this phenomenon by studying the distribution of singular values in the weight matrices of source models. They conjecture that the effectiveness of neuron alignment techniques is predicated on their ability to preferentially align the singular directions corresponding to dominant singular values. They find that the proportion of dominant singular values decreases substantially as model width increases, suggesting that wider models are easier to align. Nguyen et al. (97) show empirical results suggesting that over-parameterized models propagate dominant principal components of their hidden representations at each layer. These results support the idea that a few dominant feature directions arise in over-parameterized models. Aghajanyan et al. (2) measure the intrinsic dimensionality of language models and similarly conclude that LLMs form more compressed representations as they scale in size. These works suggest that larger models, despite their increased expressiveness, have fewer dominant directions in feature space.
@@
Some work has also attempted to clarify why over-parameterized models may display more amicable optimization behavior than smaller models. Huang et al. (54) and Li et al. (74) visualize the loss landscapes of models at various sizes and find that overparameterized models often converge to large, flat basins correlated with enhanced generalization capability (62; 54; 17; 49; 68; 29; 79) as opposed to small sharp minima. They explain this tendency by observing that such basins occupy far more volume in objective space as the dimensionality of a model increases. The increased volume of flat basins, coupled with the results presented by Simsek et al. (106) indicating that the number of permutation-equivalent objective basins corresponding to global minimizers increases rapidly as models grow in size, may explain the relative ease with which large models train: the volume of objective space corresponding to good solutions grows rapidly with model size.
@@
Figure 7. Mode Directedness: Optimal parameter settings for related tasks lie clustered together in parameter space; clusters become looser as tasks diverge in similarity. This suggests a hierarchical organization of the objective landscape with respect to semantically similar tasks. We use a handful of example natural language tasks here to illustrate this concept.
6.2.2. Pre-training Objective:
@@
Models with different pre-training objectives display different sensitivities to various merging techniques (64; 110). For example, transformer models in the T5 family (101) perform well even with only simple averaging; meanwhile, DeBERTa models (46) consistently degrade in performance when merged, even when using state-of-the-art techniques (64). This discrepancy points to an interplay between the pre-training objective and the success of model merging procedures. One possible explanation for this difference is the effect of pre-training objective on training dynamics. For example, contrastive pre-training has been observed to coerce models to lie in flatter objective basins associated with better generalization and robustness guarantees (38). Longer pre-training procedures seem to predispose models toward representations that display linear connectivity and linear relationships between activations (64; 140). Understanding this interaction is important for interpretability work that seeks to isolate task-specific behavior (92; 93; 122) or to disentangle and identify key directions in latent space (13). Some work has explored the overlap between pre-training objective, loss landscape geometry, and adversarial robustness, even demonstrating how to prevent fine-tuning pre-trained models on restricted task domains (38; 25). Further work might consider the causal effects of pre-training objective on loss landscape geometry and relate it to resulting model robustness and performance in downstream tasks.
7. Discussion
Here, we discuss how insights into loss landscape geometry from the study of model merging can be applied to enhance model interpretability and robustness.
7.1. Connections to Model Interpretability.
Techniques in interpretability research, such as activation steering, can control language style (70), reduce model toxicity (80), and edit factual associations post-training (92; 93). Such methods successfully alter model behavior via linear interventions on the activation space (121; 5), with some techniques choosing to edit weights post-training (92; 93). This largely activation-centric model of interpretability has yielded useful mechanistic interventions for controlling model behavior, however, the apparent brittleness of contemporary interpretability techniques under extensive evaluation makes clear the need for a novel perspective on model interpretability. Rigorous examinations have demonstrated that natural language explanations for neuron activation patterns display poor precision and recall while also failing to behave as expected under causal interventions (53). Even unsupervised interpretability techniques such as sparse autoencoders (13; 55; 83) have been shown to perform poorly on causal intervention benchmarks (18). In general, activation analyses—which are dependent on inference data—may not yield robust, generalizable interpretations of model components. Insights from the study of loss landscape geometry can provide researchers with tools to circumvent these pitfalls and develop more reliable interventions to guide model behavior.
Existing model merging work on task arithmetic (58), TIES-merging (130), and AdaMerge (133) demonstrates the viability of weight steering as an alternative to these activation-based techniques. The established linear correspondence between changes in model activations and changes in model weights further justifies this approach (142; 140). Weight steering has the additional advantage that it circumvents the need for extensive tuning, which is a critical downside to applying activation steering methods (112; 70; 80). For example, Barbulescu and Triantafillou (8) demonstrate how task arithmetic can be harnessed to produce superior unlearning performance in LLMs while maintaining or even improving the underlying models’ text-generation capabilities. Kadhe et al. (67) explicitly employ model merging to augment existing unlearning techniques for LLMs by splitting their unlearning dataset, unlearning each subset individually, and merging the resulting models. This technique significantly improves unlearning capabilities while maintaining performance.
@@ 7. Discussion
Future work may seek to directly apply model merging techniques to studying phenomena such as memorization (16; 103), overfitting (87), and catastrophic forgetting (69). Interpretability work in these areas can draw upon insights from the model merging literature, for instance: (i) models trained on similar tasks or exhibiting similar behaviors, can be aligned to find common subsets of parameters for a given task (3; 107; 108); (ii) models trained using distinct randomly chosen subsets of data or different hyperparameters, when merged, can help identify parameters in each source model that are relevant to the overall task invariant to the specific data or hyperparameters that were used to train each model (124; 144; 63; 108); and (iii) models with similar underlying training distributions cluster together in parameter space (44).
For example, we might choose to study how toxicity in LLMs arises by applying the methods from merging techniques such as ZipIt! (108) or OTFusion (107) to identify sets of parameters across models that are responsible for toxic behavior—potentially identifying common patterns in network connectivity. Similarly, we might attempt to associate degenerate behaviors, such as memorization, with particular features of the local loss landscape geometry, developing techniques to preemptively determine whether models are at high risk for memorizing training data. Or, we might use concepts from loss landscape geometry to determine the bases over which we can best interpret our models. For instance, Bushnaq et al. (15) explicitly invoke ideas from the study of loss landscapes to formulate a novel, more interpretable neural network parameterization. Exploring the intersection between model merging, loss landscape geometry, and interpretability is a fruitful direction for future interpretability research.
7.2. Connections to Model Robustness
Model merging techniques can aid practitioners in associating model components with semantic behaviors, such as sensitivity with respect to inputs and susceptibility to adversarial examples (e.g., robustness). Many deep learning applications that must defend against adversaries—malicious actors seeking to extract sensitive data, degrade model performance, or abuse models for inappropriate use cases—can benefit from developing novel, reliable methods to ensure model robustness. For example, in sensitive domains like healthcare, the consequences of poor model performance in diagnostic and prognostic tasks can be life-threatening—these models must be robust to adversaries (34). Healthcare providers may seek to accelerate medical research by using decentralized machine learning techniques, such as federated learning (77), to create robust models without directly sharing patient data between facilities (11). Unfortunately, decentralized model updates may risk unintentionally leaking private data by sending parameters or gradient updates across a network (27) and allowing the introduction of adversarial examples that degrade overall model performance (11; 72; 23; 136). The model merging literature can help us explain the root causes of these vulnerabilities and provide techniques to prevent them. For example, Shokri et al. (105) demonstrate that adversaries can exploit mode directedness to perform membership inference attacks on black-box models hosted in the cloud, additionally showcasing the effectiveness of mitigations such as increasing prediction entropy or applying regularization in thwarting potential adversaries. Model merging techniques such as RegMean (64) offer procedures for decentralized model merging that avoid the risk of leaking data samples. Furthermore, several recent studies in model robustness have established the effectiveness of model merging in improving existing models to be more robust and performant (38; 142; 128).
@@ 7.2. Connections to Model Robustness
Critically, the principles of mode determinism and mode directedness are implicitly or explicitly applied in various attacks on deep learning models. When attacking black-box models, it is common to train local substitute models of a similar architecture on training data for the same underlying task to develop adversarial examples (99; 10). This assumption of similarity in learned representation and behavior is encapsulated by the concept of mode directedness and is supported by the model merging and loss landscape literature. Many attacks, such as data poisoning and membership inference, can be aided by training similar model architectures on similar data distributions (105). Mode determinism implies that such attacks may be especially effective when deployed against freely available pre-trained checkpoints, as fine-tuned models stemming from the same pre-trained checkpoint will likely learn very similar representations and lie close together in the objective landscape (40; 44). This is especially a concern for fully open-source model checkpoints, as attackers can develop gradient-based attacks (i.e., developing adversarial examples) with some confidence that those attacks will work as intended against other fine-tuned models, even if they are fine-tuned on different datasets. For example, the membership inference attack proposed by Shokri et al. (105) attempts to learn whether a particular data record is present in a black-box model’s training dataset by training several “shadow models” to approximate the behavior of the target model. This attack already shows great effectiveness with only black-box access to the target model; giving adversaries access to a pre-trained version of the target model would significantly increase the attack’s severity. Similarly, Eustratiadis et al. (32) propose a gradient-based attack that can circumvent common inference-time defenses such as gradient obfuscation by estimating a smoothened version of the local loss landscape geometry around the target model. However, such gradient-based attacks might be trivially feasible given a set of open-source model weights of the same architecture that an adversary could use to develop adversarial examples. Given the popularity of the pre-train fine-tune paradigm, particularly for LLMs, there may be significant security risks associated with deploying open-source models into production scenarios.
@@
Overall, the study of model robustness is increasingly dependent on understanding loss landscape geometry (138; 128). For example, the success of gradient obfuscation against gradient-based adversaries has been linked to how gradient obfuscation alters the local loss landscape geometry to be more coarse—making it difficult to ascend for adversaries. Conversely, adversaries can effectively circumvent these defenses by navigating a smoothed surrogate loss geometry (32). Train-time defenses against adversarial examples frequently involve navigating toward robust local minima by including an adversarial robustness term in the loss—one can even thwart subsequent fine-tuning success in restricted domains by appropriately engineering the loss function (25). This has been observed to bias models toward large, flat loss basins. Indeed, a strong correlation has been established between adversarial robustness and local loss curvature, indicating that solutions in large, flat basins are more robust to adversaries (138). Some recent works even explicitly exploit mode connectivity to learn paths between vulnerable solutions that contain more robust models—outperforming fine-tuning approaches in terms of model robustness and accuracy (138). This underscores the advantages that a deep understanding of loss landscape geometry may yield in enhancing model robustness.
8. Conclusion & Future Work